Use the definition of the derivative to differentiate V=(4)/(2)\pi r^(3).


1
Expert's answer
2021-11-17T15:42:20-0500

Rewriting the expression given


"V=\frac{4}{2}\pi\>r^3"


"V=2\pi\>r^3"


Derivative is defined as "\frac{\Delta\>y}{\Delta\>x}"


In a graph of V against "r"


Derivative will be defined as "\frac{\Delta\>V}{\Delta\>r}"


Taking two general points on the graph


"(r, 2nr^3)" and "(r+\Delta\>r,2\pi[r+\Delta\>r]^3)"



"\frac{\Delta\>V}{\Delta\>r}=\frac{2\pi[r+\Delta\>r]^3-2\pi\>r^3}{(r+\Delta\>r)-r}"



"=\frac{2\pi[r^3+3r^2\Delta\>r+3r(\Delta\>r)^2+(\Delta\>r)^3-r^3]}{\Delta\>r}"



"=2\pi[3r^2+3r\Delta\>r+\Delta\>r^2]"



As "\Delta\>r\to0,"

Then "\frac{\Delta\>V}{\Delta \>r}=6\pi\>r^2"